OpenStudy (anonymous):

Write the slope-intercept form of the equation of the line described. Through (1,4), perp to y=-x-1

OpenStudy (anonymous):

perpendicular slopes are negative reciprocals. so slope needs to be 1/2 y - 4 = 1/2 (x -1 y-4 = 1/2x -1/2 y = 1/2x -1/2 + 4 y = 1/2x - 7/2